OpenStudy (anonymous):

What is the approximate distance from (2, 5) to (-2, 1)? 6 units 5 units 4 units 3 units

OpenStudy (amistre64):

4 sqrt(2) ; just like last time i believe :)

OpenStudy (anonymous):

6 units

OpenStudy (anonymous):

\[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\] \[=\sqrt{(1-5)^2+(-2-2)^2}\] \[=\sqrt{(-4)^2+(-4)^2}\] \[=\sqrt{16+16}\] \[\sqrt{32}=4\sqrt{2}\] \[=5.656\] \[=6 units\]